OpenStudy (anonymous):

what can we say about 2 non-zero vector u and v , that satisfy ||u +v|| = ||u|| + ||v|| .

OpenStudy (callisto):

||u +v|| = ||u|| + ||v|| + 2 ||u|| ||v|| cos theta, where theta = angle between the vector u and v. If u and v are both non-zero vectors, that means cos theta is 0. At what angle will cos (theta) be zero?!

OpenStudy (anonymous):

sorry , I'm not quite sure if this is what you need

OpenStudy (anonymous):

please leave those numbers alone, we don't need them

OpenStudy (callisto):

Hmm.. Why is it so complicated :S

OpenStudy (anonymous):

@Callisto , I try to figure out your formula in text book, but I cannot see. would you please tell me the name of it? or theorem or some thing about that

OpenStudy (callisto):

My bad!!! I mixed up something!! Sorry!! Ignore my comments!!

OpenStudy (anonymous):

were you writing : u dot v=|u| x|v|x cos theta?

OpenStudy (anonymous):

@ranxu6j3 I see your pic, thanks for reply but the logic is not stable.how can you get U1 +V1 become U1,U2? or I misunderstand something from yours? if so, please, correct me

OpenStudy (anonymous):

thanks for letting me know your confusion, I was taking the 2 vectors apart into verticality and parallel then I can proof the 2 sides of = are the same

OpenStudy (anonymous):

@ranxu6j3, we don't know whether \[u,v\in\mathbb{R}^2\]

OpenStudy (anonymous):

@ranxu6j3 you mean \(\vec u\) parallel to \(\vec v\)?

OpenStudy (anonymous):

@ranxu6j3 even though we have they are parallel, we must base on some thing to prove, right? some formula or concept or something else...right?

OpenStudy (anonymous):

do we? sorry , I thought with para and vert we can proof that?

OpenStudy (anonymous):

sorry I didn't see your last reply, I didn't mean v and u are para

OpenStudy (anonymous):

@SithsAndGiggles , can you explain to me what does R^2 means? tks

OpenStudy (anonymous):

@ranxu6j3 the proof must be strong and logic, not intuitive or "trivially"

OpenStudy (anonymous):

Triangle inequality? We know that, in general, \[||u+v||\le ||u||+||v||.\] I think some manipulation of this will help. @ranxu6j3, \[\mathbb{R}^2=\mathbb{R}\times\mathbb{R}, \text{ defined by }\\ \mathbb{R}^2:=\left\{(x,y):x,y\in\mathbb{R}\right\}\] (the x stands for Cartesian product)

OpenStudy (anonymous):

sorry I am not a native speaker so that I might misunderstand you guys when Math in English. I apologize for making you confused, but i'll try. could you tell me that should we try to proof the "="(since I am not sure what exactly you want)

OpenStudy (anonymous):

thanks @SithsAndGiggles , i got your meaning

OpenStudy (anonymous):

My point is that, although your logic/intuition may work for R², it may not necessarily work for higher dimensions or other vector spaces.

OpenStudy (anonymous):

@ranxu6j3 means = (equal) @SithsAndGiggles that is inequality formula, but it doesn't help because it has no condition to get extremely = or extremely <

OpenStudy (anonymous):

yeah, i got it , @SithsAndGiggles , so do we have any limitation or condition in this case, @Hoa ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

does the case says we are in 2d ?

OpenStudy (anonymous):

I post exactly what the book print

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

oh that's why my pic doesn't work

OpenStudy (anonymous):

Okay, I'm pretty sure I know what the answer is, but I'm not sure how to formally obtain it through a proof...

OpenStudy (anonymous):

@SithsAndGiggles write it out, others can contribute

OpenStudy (anonymous):

If u=v, then the equality holds:\[||u+v||=||u+u||=2||u||=||u||+||u||\]

OpenStudy (anonymous):

I'll see if I can write something up...

OpenStudy (anonymous):

I think it's just a fact that comes from the triangle inequality; the equality holds when the given vectors are equal, and doesn't when they're not.

OpenStudy (anonymous):

everybody, what do you think?

OpenStudy (anonymous):

my reply was met a problem , if , for example : u=(1,2) v=(1,-1) , then those i thought would be broken , and i am not quite sure how do we proof it with triangle inequality

OpenStudy (anonymous):

The fact does come from the triangle inequality. However, the vectors do not need to be equal. They just need to be linearly dependent (or collinear). I think I have a proof for it, give me a minute to check it though.

OpenStudy (anonymous):

i have an idea, everybody check whether it is right or not\[||\vec u +\vec v|| = \sqrt{(a+d)^2 +(b+e)^2 +(c+f)^2}where (\vec u = (a,b,c) and \vec v =(d,e,f)\]

OpenStudy (anonymous):

@Hoa, if you are right, you'd only be making the case for R³, and not for a general vector space.

OpenStudy (anonymous):

||\(\vec u\)||+||\(\vec v\)|| = \[\sqrt{a^2 + b^2 +c^2 } + \sqrt{d^2 +e^2 +f^2}\]

OpenStudy (anonymous):

i think this is the same as what i wrote on my pic, but this time in 3D

OpenStudy (anonymous):

make them equal and find out the condition to get that become true

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

I understand what you wrote, but it must be mathematically proved, right?

OpenStudy (anonymous):

yeah so it's good that you make it into math tks, and I think what you wrote can be used

OpenStudy (anonymous):

i can finally get back here, but i can't see those you post still probs, sorry for not continuing, i'll try later @hoa

OpenStudy (callisto):

||u +v|| = ||u|| + ||v|| Square both sides ||u +v||^2 = (||u|| + ||v|| )^2 ||u +v||^2 = ||u||^2 + ||v||^2 + 2 ||u|| ||v|| Consider the left side ||u +v||^2 = \((u+v)\cdot (u+v)\) = \(||u||^2 +||v||^2 + 2u\cdot v\) = ||u||^2 + ||v||^2 + 2 ||u|| ||v|| cos theta So, we get \(||u||^2 + ||v||^2 + 2 ||u||\ ||v|| cos \theta = ||u||^2 + ||v||^2 + 2 ||u|| \ ||v||\) Cancel the common terms: \(2 ||u||\ ||v|| cos \theta =2 ||u|| \ ||v||\) Cancel the common factors: \[cos \theta =1\] Wow!! This tells you something, doesn't it?

OpenStudy (anonymous):

by using geometry , I have 2 results, \(\vec u\)parallel \(\vec v\) either same direction and opposite , I still have the same result since they are ||...|| , is it right?

OpenStudy (anonymous):

I draw 3 circles with the same radius and let them cut each other, like

OpenStudy (callisto):

If the direction doesn't matter, why not just write "u and v are parallel to each other"??

OpenStudy (anonymous):

but if they don't have the same initial point, how the result comes true? I mean if they // but overlap a while

OpenStudy (anonymous):

oh oh, my bad, you are right

OpenStudy (anonymous):

the length. thanks a lot

OpenStudy (callisto):

Welcome :) Sorry for my mistakes earlier!!

OpenStudy (anonymous):

@Callisto thanks for your idea.